OpenStudy (anonymous):

For the system shown below, what are the coordinates of the solution that lies in quadrant I? 3y^2-5x^2=7 3y^2-x^2=23

OpenStudy (anonymous):

(-2,3)??

OpenStudy (anonymous):

or (2,3)?

hero (hero):

Hint: 1.) Isolate 3y^2: 3y^2 = 7 + 5x^2 3y^2 = 23 + x^2 2.) Set 3y^2 = 3y^2 7 + 5x^2 = 23 + x^2 3.) Finish the rest

OpenStudy (anonymous):

I thought that there are 4 solutions, one in each quadrant: (± 2, ± 3) Q I : (2, 3) Q II : (−2, 3) Q III : (−2, −3) Q IV : (2, −3)

hero (hero):

Yeah, you're right. Four solutions

OpenStudy (anonymous):

so (2,3) is my answer?

hero (hero):

You have four solutions, not one.

hero (hero):

Oh, yeah, for quadrant 1, (2,3)

OpenStudy (anonymous):

Cool thanks soo much :)